In ∆ABC, a =3 ,b =8, and sin A =5\13, then find the number of triangles

Grade:11

1 Answers

Sujay Yadav
6 years ago
SinA/a =sin B/b. 5/13|3=sin B/8. =15/13=sinB/8. =sin B =120/13>1this is not possible =0 so this is the answer of this question the value of sin B is greater then 1 so answer is 0